A saturated solution of lead (II) iodide, PbI2 has an iodide concentration of 3.0 × 10-3 mol/ L. Calculate the solubility constant, Ksp, for lead (II) iodide. PbI2 (s) ß---à Pb2+(aq) + 2I-(aq)A saturated solution of lead (II) iodide, Pbl2 has an iodide concentration of 3.0 x 10-3 mol/ L. Calculate the solubility constant, Ksp, for lead (II) iodide Pbl2 (s) 8 3.5 x 10-8 5.0 x 10-8 2.8 x 10-8 1.4 x 10-8 -à Pb2 (aq)+2I(aq)
Answer #1

At equilibrium:

PbI2 <----> Pb2+ + 2 I-

   s 2s

Given:

2s = 3*10^-3 M

So,

s = 1.5*10^-3 M

use:

Ksp = [Pb2+][I-]^2

Ksp = (s)*(2s)^2

Ksp = 4(s)^3

Ksp = 4(1.5*10^-3)^3

Ksp = 1.4*10^-8

Answer: 1.4*10^-8